OpenStudy (anonymous):

there is a right triangle. the hypotinuse lenghth is 13 the opposite side length is 12 what is the adjacent side lenth

OpenStudy (anonymous):

5

OpenStudy (anonymous):

169 - 144 = 25 = 5^2

OpenStudy (anonymous):

do you guys knew about tan, sin, cos i have to use that to get the answer

OpenStudy (anonymous):

(13)^2 = (12)^2 + x^2 169 = 144 +x^2 169-144 = x^2 25 = x^2 5 =x

OpenStudy (anonymous):

is that using one of those three things

OpenStudy (anonymous):

nope its simply using pythagoras theorem

OpenStudy (anonymous):

Look, \[\sin(x) = 12/13\]

OpenStudy (anonymous):

the adjacent side is \[13 * \cos(x) \]

OpenStudy (anonymous):

and pythagoras gives \[ \sin^2(x) + \cos^2(x) = 1 \]

OpenStudy (anonymous):

why to include cos and sin when we can directly get the answer?

OpenStudy (anonymous):

so i guess you could say that, but its really unnnecessary

OpenStudy (anonymous):

my teacher showed us these three things today so we have to use t for our hw

OpenStudy (anonymous):

There is no need for the vocabulary of trigonometry to enter this problem.

OpenStudy (anonymous):

oo really